Problem 3.10. How many ways are there to seat 8 people, consisting of 4 couples, in a row of seats (8 seats wide) if all couples are to get adjacent seat?
Answer by jim_thompson5910(35256)

Each couple combines to form a single unit. So you really have 8/2 = 4 units. There are 4! = 4*3*2*1 = 24 ways to arrange these units. Now because you can arrange the people of each individual unit (focusing on just one specific unit), there are really 2*24 = 48 ways to fully arrange the people so they are adjacent to their partner.